[figure: 9.7 cm]

THE ELECTRICAL CONDUCTIVITY IN THE OFF-STOICHIOMETRIC HEUSLER Fe_{2}VX (X=Al, Ga)

J. Deniszczyk^{a} and W. Borgieł^{b} a Institute of Physics and Chemistry of Metals, University of Silesia 40-007 Katowice, Bankowa 12, Poland b Chełkowski Institute of Physics, University of Silesia 40-007 Katowice, Uniwersytecka 4, Poland

The Fe_{2}VX (X=Al, Ga) Heusler compounds exhibit several properties uncommon among the 3d-transition metal intermetallic compounds. The most intriguing is the semiconductor-like behavior of the electrical resistivity with simultaneous anomalously large value of electronic specific coefficient γ. The calculated electronic structure of the Fe_{2}VAl characterizes a deep, 0.5 eV wide, pseudo-gap centered at Fermi level (ε_{F}). Our recent electronic structure calculations of the off-stoichiometric Fe_{(2+y)}V_{(1−y)}X (y = ±0.06) compositions have proved the existence of the narrow band located just at ε_{F}. In the non-magnetic state of the y > 0 composition the band appeared to be the almost fully occupied upper Hubbard band of the d-e_{g} states of the impurity Fe atoms. The presence of the band explains the large value of the γ coefficient. It may also response for the negative value of the temperature coefficient of the DC-resistivity. To investigate this we considered the two-band model of the strongly correlated narrow band coupled by the hybridization interaction with the quasi-free electron band. The model was solved by the CPA approach within the Alloy Analogy approximation and the Kubo-Greenwood formula for the DC-conductivity was applied. The calculated temperature dependence of the DC-resistivity qualitatively agree with the measured one.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : J. Deniszczyk

Address for correspondence : Institute of Physics and Chemistry of Metals University of Silesia 40-007 Katowice, Bankowa 12

Email address : jdeni@us.edu.pl